# SMITIN: Self-Monitored Inference-Time INtervention for Generative Music Transformers

### v1

@@ 3.2 MusicGen Probing
3.3 Qualitative Evaluation
We qualitatively analyze the performance of top-$K$ probes’ monitoring capability through Fig. 5. The red dashed line in each plot indicates the threshold value $\tau$, as mentioned in Section 2.5. From the left side of the figure, it is apparent that the median (yellow line) prediction more closely aligns with the actual presence of the target instrument than the mean prediction (green line). Consequently, we use $\bar{\mathcal{C}}(t)=\text{med}(\mathcal{C}(t))$ to monitor during inference time.
The sub-plots on the right illustrate how the hyper-parameter $s$ influences the outcome of the audio continuation task. We observe a relationship between the intervention frequency $s$ and the rapidity with which the model aligns toward the target factor. This suggests that more frequent interventions cause the transformer to prioritize the target factor more heavily in subsequent token generation, potentially at the expense of referencing previous musical context. We discuss this trade-off between intervention quantity and musical coherence throughout this section.
3.4 Objective Evaluation
Utilizing the metrics introduced in Section 3.1.2, we objectively evaluate our methods on Audio Continuation and Text-to-Music. We compare with baseline methods that do not incorporate ITI, such as MusicGen’s standard unconditional and text-prompt conditioned generations. Additionally, we benchmark against the original ITI method, which applies a uniform intervention at every time step (i.e., $s=1$) with a configuration of $\alpha=5.0$ and $K=64$ to apply the intervention with the equal amount $\alpha$ to each of the $K$ selected heads. For Audio Continuation, we examine another ITI approach, weight decay, where $\alpha$ is progressively reduced across time steps during the generation process. For Text-to-Music, our evaluation includes both our default configuration ($\alpha=5.0$) and a stronger intervention level of $\alpha=10.0$ to observe the trade-off between the intervention strength and the preservation of musical fluency within the generated samples.
Table 3.2 presents the objective evaluation for the Audio Continuation task. The baseline methods reveal that MusicGen inherently tends to add certain instruments without explicit direction with varying success rates across instruments; drums and bass are more easily added compared to guitar and piano. Intriguingly, text-prompted generation does not consistently achieve the target instruction. Intervention methods, on the other hand, consistently outperform baseline approaches in successfully directing MusicGen toward adding the desired instruments, as indicated by the success rate. The original ITI, applying a constant level of intervention at each generation step, leads to a higher success rate but significantly alters the distribution of generated content, which can be observed via the similarity score for maintaining musical consistency. In contrast, the weight decay ITI approach has a subtler influence on the quality but produces a lower success rate. Our proposed SMITIN shows a notable 10.5% jump over text-prompt conditioning, and is better at retaining the model’s output distribution and generating consistent music. When SMITIN is deployed in conjunction with text-conditioning, it outperforms text-based instructions, but not to the same extent as when used independently. This result opens up the potential for integrating ITI with textual directives, allowing a dual axis of control—semantic guidance via text and fine-tuned adjustments through ITI.
For the Text-to-Music task, we explore the interplay of ITI with text prompts, as shown in Table 3.2. In this context, we utilize MusicCaps text prompts as the basis of MusicGen’s performance. Consistent with earlier observations, the original ITI method, despite showing the highest success rate in instrument addition, performs the lowest in terms of FAD, indicating a more pronounced shift from the natural music distribution. SMITIN, in its base configuration ($\alpha=5.0$), achieves an average performance on par with additional text guidance. We further analyze SMITIN with an increased intervention strength of $\alpha=10.0$, and observe that a clear trade-off emerges between the success rate and the naturalness of the generated music. This outcome highlights the flexibility of SMITIN, offering users a tunable “knob” to balance between precision in achieving specific musical characteristics and maintaining the authenticity of the musical piece.
3.5 Subjective Evaluation
Table 3: Average and 95% confidence interval of subjective listening tests for overall audio quality across four instrument addition interventions for audio continuation and text-to-music experiments.
@@ 3.5 Subjective Evaluation
To further validate the effectiveness of our intervention techniques on MusicGen, we performed a subjective listening test through Amazon Mechanical Turk following the best practices from (Ribeiro et al., 2011). The purpose of this test was not to evaluate the quality and relevance of MusicGen as this was already explored in (Copet et al., 2023), but rather to provide quantitative evidence that: (1) our proposed success rate objective metric correlates with human perception, and (2) our proposed intervention techniques are not detrimental to overall audio quality. To do this, we assessed whether human listeners could detect the presence of the instruments added by our intervention technique, and asked them to rate samples in terms of overall quality on a scale of 1-5. We also included real music samples from MoisesDB to provide a performance baseline. Table 3 compares the overall audio quality ratings, where we note that our proposed intervention techniques perform comparably to applying no intervention. Furthermore, we computed the Spearman correlation coefficient between the objective success rate metric and the average score across human raters asked to detect the presence of the instrument added by the intervention techniques. Across all intervention algorithms, we achieved $\rho=0.48$ for audio continuation and $\rho=0.67$ for text-to-music, both with p-values much less than 0.05, indicating that our proposed success rate metric correlates with the ability of humans to detect instruments added through intervention. Further results and details on the experimental setup are provided in Appendix A.
@@
4 Analysis
@@ 4.1 Intervention with Multiple Directions
We explore the application of ITI in scenarios where multiple musical elements are introduced simultaneously. In our setup, we focus on the continuation of music that has one instrument present and assess the addition of three others. For instance, the intervention task is to generate drums, bass, and piano when the input music has only guitar. We consider two types of success rates, an individual one for each instrument as before, and a simultaneous one to assess how often all 3 target instruments were added together. We consider a successful case for a song if each instrument is deemed present more than 50 % of the time.
The summarized results in Table 4 reveal that while text prompts yield higher success rates for individual instruments, their generated output diverges from the input music’s characteristics. This indicates that text conditioning may not adequately consider the input and opts to generate new content based solely on the given instruction. Without ITI, MusicGen tends to produce continuations focused on a single instrument, which deviates from the desired complex mixtures, as evidenced by the FAD scores. SMITIN, however, significantly outperforms text prompts in preserving input music similarity, and further enables more granular control over each musical aspect, leading to a higher success rate of generating all desired instruments in unison. This fine-grained control mechanism bolsters SMITIN’s potential as a robust tool for complex music generation tasks where maintaining the essence of the input is crucial.
@@
4.2 Generating “Realistic” Music
To generate more realistic music through ITI, we fit probes to discern between real music (DISCO-10M dataset’s DISCO-200K-high-quality subset (Lanzendörfer et al., 2023)) and synthetic outputs from MusicGen itself. Surprisingly, the probing showed high performance with the best-performing probe achieving a 96.2% accuracy rate and an average accuracy of 77.5% in distinguishing real from synthetic music (See probing results in Appendix LABEL:fig:probe_acc_real_fake). Leveraging this result, we then steer MusicGen in an audio continuation task, using real-world music inputs from the DISCO-10M dataset to generate continuations that maintain the “realistic” quality.
Table 6: Ablation study on the number $n$ of paired data for probing (performance is averaged over all instruments, shows full-set success rate).
$n$ = 10 $31.0$ $0.357$ $0.909$ 79.1
$n$ = 100 $31.8$ $0.349$ $0.914$ 82.6
$n$ = 500 $27.3$ $0.341$ $0.915$ 84.4
$n$ = 1000 $25.5$ $0.340$ $0.914$ 84.8
[…]
4.3 Effects on Number of Probing Data
To investigate the influence of probing data quantity on SMITIN, we test the system with varying sizes of probing datasets randomly selected from the complete set. These subsets consist of $n=10$, $100$, $500$, and $1000$ data pairs, equivalent to music durations of $1$, $10$, $50$, and $100$ minutes respectively. The results are presented in Table 4.2, where we show the “full-set success rate” which is computed using the classifier probes trained with the full dataset every time.
The results indicate that the system demonstrates reasonable results even with a minimal dataset of just 1 minute. As the data size increases to $n=500$, the objective metrics stabilize, indicating an optimal data quantity threshold for effective probe training. An interesting observation from our results is the higher full-set success rates associated with smaller datasets. This trend suggests that probes trained on limited data may have lower confidence in their decisions, prompting them to recommend more interventions during the ITI process to meet the success threshold. At the same time, we observe a clear correlation between better probes (represented here through the accuracy of the best probe) and better ITI performance, whether measured as higher full-set success rate or similarity, or lower FAD.
This altogether has practical implications for the application of ITI in real-world scenarios. It further supports the extensive use of probing as the basis for effective ITI design. But it also suggests that users can effectively leverage ITI with only a small amount of their data to control generation. It means that effective ITI can be accessible and achievable without the need for extensive datasets, which is particularly beneficial for individual artists or smaller studios. This accessibility broadens the potential for creative and personalized applications of ITI in music generation, making it a versatile tool for a wide range of users.

### v2

@@ 4.3 Objective Evaluation
Table 4.3 presents the objective evaluation for the audio continuation task. The baseline methods reveal that MusicGen inherently tends to add certain instruments without explicit direction with varying success rates across instruments; drums and bass are more easily added compared to guitar and piano. Intriguingly, text-prompted generation does not consistently achieve the target instruction. Intervention methods, on the other hand, consistently outperform baseline approaches in successfully directing MusicGen toward adding the desired instruments, as indicated by the success rate. The original ITI, applying a constant level of intervention at each generation step, leads to a higher success rate but significantly alters the distribution of generated content, which can be observed via the similarity score for maintaining musical consistency. In contrast, the weight decay ITI approach has a subtler influence on the quality but produces a lower success rate. Our proposed SMITIN shows a notable 10.5% jump over text-prompt conditioning and is better at retaining the model’s output distribution and generating consistent music. When SMITIN is deployed in conjunction with text conditioning, it outperforms text-based instructions, but not to the same extent as when used independently (likely because we do not intervene on the text cross-attention layers).
@@
For the text-to-music task, we explore the interplay of ITI with text prompts, as shown in Table 4.3. In this context, we utilize MusicCaps text prompts as the basis of MusicGen’s performance. Consistent with earlier observations, the original ITI method, despite showing the highest success rate in instrument addition, performs the lowest in terms of FAD, indicating a more pronounced shift from the natural music distribution. SMITIN, in its base configuration ($\alpha=5.0$), achieves an average performance on par with additional text guidance. We further analyze SMITIN with an increased intervention strength of $\alpha=10.0$ and observe that a clear trade-off emerges between the success rate and the naturalness of the generated music. This outcome highlights the flexibility of SMITIN, offering users a tunable “knob” to balance between precision in achieving specific musical characteristics and maintaining the authenticity of the musical piece.
Table 4: Average and 95% confidence interval of subjective listening tests for overall audio quality across four instrument addition interventions for audio continuation and text-to-music.
@@
Table 5: Subjective results across multiple algorithms, comparing human (Hum.) detection on the presence of instruments added by intervention, with Success Rate (Alg.). All Spearman correlation coefficients ($\rho$) have p-values less than 1e-5.
Audio continuation Text-to-music
Top Bottom Top Bottom
Inst. $\rho$ Hum. Alg. Hum. Alg. $\rho$ Hum. Alg. Hum. Alg.
drums $0.76$ $0.88$ $0.81$ $0.19$ $0.13$ $0.80$ $0.95$ $0.95$ $0.16$ $0.18$
bass $0.42$ $0.78$ $0.89$ $0.48$ $0.06$ $0.58$ $0.80$ $0.96$ $0.38$ $0.13$
guitar $0.40$ $0.78$ $0.76$ $0.52$ $0.15$ $0.67$ $0.86$ $0.88$ $0.34$ $0.14$
piano $0.32$ $0.60$ $0.72$ $0.33$ $0.32$ $0.55$ $0.70$ $0.79$ $0.31$ $0.23$
avg. $0.48$ $0.76$ $0.80$ $0.39$ $0.17$ $0.67$ $0.83$ $0.90$ $0.30$ $0.17$
Table 6: Spearman correlation of instrument recognition between human decisions and recognition models. All correlations have p-values less than 1e-5.
Correlation Coefficient $\rho$ ($\uparrow$)
Method drums bass guitar piano avg.
Success Rate 0.78 0.48 0.57 0.45 0.58
YAMNet 0.80 0.37 0.56 0.33 0.57
PaSST 0.79 0.49 0.59 0.43 0.57
CLAP 0.80 0.53 0.56 0.52 0.65
ConvNeXt 0.69 0.52 0.51 0.46 0.57
4.4 Subjective Evaluation
To further validate our intervention techniques on MusicGen, we performed a subjective listening test through Amazon Mechanical Turk following the best practices from [46]. The purpose of this test was not to evaluate the quality and relevance of MusicGen as this was already explored in [2], but rather to provide evidence that: (1) our proposed success rate objective metric correlates with human perception, and (2) our proposed intervention techniques are not detrimental to overall audio quality. To do this, we assessed whether humans could detect the presence of the instruments added by our intervention technique, and asked them to rate samples in terms of overall quality on a scale of 1-5. We also included real music samples from MoisesDB to provide a performance baseline. In selecting files for each algorithm and instrument, we randomly selected 20 files above the median objective success rate and 20 files below the median. This leads to 1280 generated files to be rated (20 files x 2 top/bottom x 4 instruments x 4 algorithms x 2 audio continuation/text-to-music). All files in the listening test were ten seconds long and normalized to -12 loudness units full scale (LUFS) [47]. We obtain at least 3 ratings per file. For the audio continuation experiments, we ask raters to ignore the first 3 seconds of each audio file, as that conditioning signal intentionally does not include the target instrument.
Table 4 displays the overall objective quality results. In general, the performance of all algorithms is quite similar, with SMITIN and no intervention having the best overall quality. In addition to overall quality, we also ask whether listeners can detect the presence of the instruments targeted by our intervention. This serves to help validate our success rate objective metric. We also compare the average human score with the average success rate objective metric (Alg.) for the top-ranked and bottom-ranked files selected for the listening test on each generation task in Table 5. For the top-ranked files the scores appear to match quite well, while for the bottom-ranked files, it seems human raters tend to overestimate the presence of most instruments compared to the objective metric. We also hypothesize that the lower match for audio continuation compared to text-to-music may be due in part to the test being more subjectively difficult as raters have to focus on the end of the audio file under test.
Table 6 compares the Spearman correlation coefficient $\rho$ computed between each of the instrument recognition systems discussed in Section 4.1.2) and the average human rating computed across all processing algorithms. Notably, using CLAP embeddings yielded the highest $\rho$ value with human decisions, while the other methods produced similar values. This result further supports the validity of using the Success Rate on generated samples. Not only does the Success Rate show a strong correlation with other recognition models, but the high correlation observed between the Success Rate and CLAP in Table 1 strengthens its reliability as an objective metric for controlling and evaluating generative music.
5 Ablations and Practical Considerations
@@
Intervention with Multiple Directions. We explore ITI in scenarios where multiple musical elements are introduced simultaneously. We focus on the continuation of music that has one instrument present and assesses the addition of three others. For instance, the intervention task generates a continuation with drums, bass, and piano when the input music only has guitar. We consider two types of success rates, an individual one for each instrument as before, and a simultaneous one to assess how often all 3 target instruments were added together. We consider a successful case for a song if each instrument is deemed present more than 50% of the time.
The summarized results in Table 7 reveal that while text prompts yield higher success rates for individual instruments, their generated output diverges from the input music’s characteristics. This indicates that text conditioning may not adequately consider the input and opts to generate new content based solely on the given instruction. Without ITI, MusicGen tends to produce continuations focused on a single instrument, which deviates from the desired complex mixtures, as evidenced by the FAD scores. SMITIN, however, significantly outperforms text prompts in preserving input music similarity, and further enables more granular control over each musical aspect, leading to a higher success rate of jointly generating all desired instruments. This fine-grained control mechanism bolsters SMITIN’s potential as a robust tool for complex music generation tasks where maintaining the essence of the input is crucial.
Generating “Realistic” Music. To generate more realistic music through ITI, we fit probes to discern between real music (DISCO-10M dataset’s DISCO-200K-high-quality subset [48]) and synthetic outputs from MusicGen itself. Despite MusicGen being trained only on real music, the probing showed high performance with the best-performing probe achieving a 96.2% accuracy rate and an average accuracy of 77.5% in distinguishing real from synthetic music. Leveraging this result, we steer MusicGen in an audio continuation task, using real-world music inputs from the DISCO-10M dataset to generate continuations that maintain the “realistic” quality.
@@
Table 8: Ablation on amount of probing data (performance is averaged over all instruments, shows full-set success rate).
$n$ = 10 $31.0$ $0.357$ $0.909$ 79.1
$n$ = 100 $31.8$ $0.349$ $0.914$ 82.6
$n$ = 500 $27.3$ $0.341$ $0.915$ 84.4
$n$ = 1000 $25.5$ $0.340$ $0.914$ 84.8
[…]
Effects of Number of Probing Data. To investigate the influence of probing data quantity on SMITIN, we test the system with varying sizes of probing datasets randomly selected from the complete set. These subsets consist of $n=10$, $100$, $500$, and $1000$ data pairs, equivalent to music durations of $1$, $10$, $50$, and $100$ minutes respectively. The results are presented in Table 5, where we show the “full-set success rate” which is computed using the classifier probes trained with the full dataset.
The results indicate that the system demonstrates reasonable results even with a minimal dataset of just 1 minute. As the data size increases to $n=500$ (50 minutes), the objective metrics stabilize, indicating an optimal data quantity threshold for effective probe training. An interesting observation from our results is the higher full-set success rates associated with smaller datasets. This trend suggests that probes trained on limited data may have lower confidence in their decisions, prompting them to recommend more interventions during the ITI process to meet the success threshold. At the same time, we observe a clear correlation between better probes (represented here through the accuracy of the best probe) and better ITI performance, whether measured as a higher full-set success rate or similarity, or lower FAD.
This suggests that users can effectively leverage ITI with only a small amount of data to control generation. It means that effective ITI for controlling any musical trait can be accessible and achievable without the need for extensive datasets, which is particularly beneficial for individual artists or smaller studios. This broadens the potential for creative and personalized applications of ITI in music generation, making it a versatile tool for a wide range of users.
6 Conclusions and Future Work
We proposed a novel approach for inference-time control of generative music transformers, which self-monitors probe accuracy to impose desired musical traits while maintaining overall music quality. A limitation of our approach is the dependence of performance on probe accuracy, that is, if the pre-trained transformer has not accurately learned a concept, or if the probe training set is inadequate, the success rate of our intervention technique may suffer (although it still could be creatively useful). In the future, we plan to explore different generative model architectures [5, 21, 49] and interventions using musical traits such as genre, emotion, etc. Furthermore, we hope to investigate building “knobs” for interactively controlling the hyper-parameters necessary for inference-time control to enable new technical and creative interactions.
Appendix
6.1 Analytical Insights from Probing
Table 9: Comparison of multi-label instrument recognition for head-wise probing on MusicGen and a supervised model.
Accuracy / F1 Score Num.
Model vocals bass drums other avg. Param.
[…]
Following the probing method outlined in the main paper, we investigate additional downstream tasks through probing to present more objective results and enhance our understanding. First, we assess MusicGen’s probing capabilities by comparing them to ConvNeXt [39], a model trained via supervised learning. Subsequently, we contrast our results with prior work on music probing [31] which uses entire attention layer outputs from Jukebox, another generative music transformer model.
6.1.1 Multi-label Instrument Recognition
Following up on the performance of single-instrument recognition, we compare the capabilities of individual attention heads in MusicGen against a supervised model trained for the multi-label instrument recognition task, i.e, identifying the presence of all instruments of a given music clip. This is achieved by fitting multiple logistic regression classifiers (probes) for each instrument class using again the extracted intermediate MusicGen activations at the last time step. For the final evaluation, we select the best-performing probe for each instrument class, noting that these may originate from different attention layers ($l$) and heads ($h$). Following the methodology described in [50], we conduct recognition tasks on 3-second music segments from the MUSDB dataset [33]. We also report the performance of a model trained in a supervised manner specifically for instrument recognition [50], utilizing the ‘tiny’ configuration of ConvNeXt [39].
The objective results for multi-label instrument recognition are detailed in Table 9. We find that the performance of MusicGen’s probes is comparable to that of the supervised method. However, across all configurations of MusicGen, we note strong performance in terms of the F1 score. Specifically, MusicGen exhibits exceptional performance in recognizing drums, whereas its least effective performance is observed in the other category. This discrepancy likely arises from the ambiguous definition of the other category in the MUSDB18 dataset, with the supervised method gaining an advantage by explicitly training on the dataset’s label, thereby enhancing its ability to identify other.
Table 10: Performance comparison of layer-wise and head-wise probing on generative music transformers. For music tagging, we report the performance of the best-performing probe across all classes, with the ensemble result of the best-performing probes for each class indicated within parentheses.
Dataset MTT GTZAN GS EMO
Task Tagging Genre Key Emotion Num. Param.
[…]
Jukebox $91.5$ $41.4$ $79.7$ $66.7$ $72.1$ $61.7$ $69.9$ 4.8K 5B
[…]
6.1.2 Various Music Downstream Tasks
To expand our understanding of MusicGen’s capabilities beyond instrument recognition, we conducted probing on a set of more general music understanding tasks. These include music tagging (MTT) [51], genre classification (GTZAN) [52], key detection (GS) [53], and emotion recognition (EMO) [54]. We adopt appropriate regression models for probing for these tasks: multinomial logistic regression for multi-class classification tasks, multiple logistic regressions for multi-label classification tasks, and linear regression for regression tasks. For these tasks, we set the input duration as 29.0 seconds, and let the probes take the last time step activation output for classification.
Table 10 presents the probing results for various configurations of MusicGen using our attention head probing, and a prior music probing work [31] that relies on a comparative model, Jukebox [1], but uses a different probing methodology:
  1. 1.
    [31] probed Jukebox using entire intermediate attention layers, whereas our approach evaluates performance based on individual attention heads.
  2. 2.
    [31] employed a one-layer MLP with 512 hidden units. In preliminary experiments, using a one-layer MLP for head-wise probes showed limited benefits, possibly because the attention dimension per head that we use is 75 times smaller than that of Jukebox.
Jukebox probing typically shows superior performance, maybe due to the model’s larger number of parameters and activation dimension. In music tagging, we present both the performance of the best-performing probe across all classes and the ensemble result of the best-performing probes for each class, indicated within parentheses. We observe an improved performance with the ensemble technique, though the improvement is typically not substantial. Interestingly, this means that we can find a single head capable of understanding multiple attributes simultaneously, a notable feature given the MTT dataset comprises 188 different tags. While a larger number of parameters generally correlates with enhanced performance, the melody configuration outperforms others in key detection. This superior performance is likely attributed to the model’s training with a melody conditioning objective [2].
[…]
6.2 Subjective Listening Test Setup and Detailed Results
In selecting files for the listening test discussed in the main paper, for each algorithm and instrument addition intervention, we randomly select 20 files above the median objective success rate, and 20 files below the median. This leads to 1280 generated files to be rated (20 files x 2 top/bottom x 4 instruments x 4 algorithms x 2 audio continuation/text-to music). We additionally include real music samples from MoisesDB as references. All audio files included in the listening test were ten seconds in length and normalized to -12 loudness units full scale (LUFS) [47]. We obtain at least 3 ratings per file. For the audio continuation experiments, we ask raters to ignore the first 3 seconds of each audio file, as that conditioning signal intentionally does not include the target instrument. In addition to following the best practices from [46], we limited participation to “Mechanical Turk Masters” for which Amazon charged a \$0.005 fee per assignment, and paid \$0.10 per rated file.
In addition to overall quality, we also ask whether listeners can detect the presence of the instruments targeted by our intervention. This serves to help validate our success rate objective metric, which was shown and discussed in the main paper.
Audio continuation Text-to-music
Method drums bass guitar piano avg drums bass guitar piano avg
MoisesDB 3.98 $\pm$ 0.29
Table 12: Intervention with Multiple Directions (Individual Success Rate).
Individual Success Rate [%]
start with drums start with bass start with guitar start with piano
Table 13: Intervention with Multiple Directions (Simultaneous Success Rate, FAD, and Similarity).
Table 14: Objective evaluation on Instrument Removal.
[…]
@@ 6.3 Intervention with Multiple Directions
We explore the application of ITI in scenarios where multiple musical elements are introduced simultaneously. In our setup, we focus on the continuation of music that has one instrument present and assesses the addition of three others. For instance, the intervention task is to generate a continuation with drums, bass, and piano when the input music only has guitar. We consider two types of success rates, an individual one for each instrument as before, and a simultaneous one to assess how often all 3 target instruments were added together. We consider a successful case for a song if each instrument is deemed present more than 50% of the time.
The summarized results in Tables 12 and 13 reveal that while text prompts yield higher success rates for individual instruments, their generated output diverges from the input music’s characteristics. This indicates that text conditioning may not adequately consider the input and opts to generate new content based solely on the given instruction. Without ITI, MusicGen tends to produce continuations focused on a single instrument, which deviates from the desired complex mixtures, as evidenced by the FAD scores. SMITIN, however, significantly outperforms text prompts in preserving input music similarity, and further enables more granular control over each musical aspect, leading to a higher success rate of jointly generating all desired instruments. This fine-grained control mechanism bolsters SMITIN’s potential as a robust tool for complex music generation tasks where maintaining the essence of the input is crucial.
6.4 Instrument Removal
We explore an additional application: removing the target instrument during audio continuation using the same probes employed for instrument addition. To accomplish this, we apply negative intervention strength $\alpha$ to eliminate the target instrument from the given input music. The hyperparameter settings for this task are configured as $\alpha=-10.0$ for both original ITI and SMITIN, with $s=1$ and $\tau=0.5$ for SMITIN. The success rate in this context is defined as $\bar{\mathcal{C}}(t)<\tau$, where $\tau=0.5$.
Table 6.2 reveals an inverse tendency for each instrument’s removal: drums and bass tend to be continuously generated, whereas guitar and piano are more likely to be omitted during the generation process. Moreover, we find that both text-conditioning prompts: “remove <inst.>”, and “no <inst.>” are ineffective in removing the target instrument. This outcome suggests another aspect of controllability that is challenging to achieve with text prompts alone.